OpenStudy (prestons2019):

An initial deposit of $50 is made into an account that had a 5% interest rate compounded annually. Which expression shows the amount of money in the account after t years?

OpenStudy (thoritegem):

Do you have any answers to choose from?

OpenStudy (prestons2019):

a=50(105)^t, b=50(1.05)^t, c=50(.05)^t, d=50(5)^t

OpenStudy (thoritegem):

Ok, so do you know what 5% of 50 is?

OpenStudy (prestons2019):

yes its 2.5

OpenStudy (1234567890):

5% of 50 is 10.

OpenStudy (prestons2019):

my bad its ten not 2.5